OpenStudy (anonymous):

-4+3y=12 slope intercept form ?

OpenStudy (hba):

Find slope. y=mx+b

OpenStudy (mrdoe):

(y-y1)=m(x-x1)

OpenStudy (mrdoe):

is the point slope form

OpenStudy (hba):

Find the slope and then use the point slope form provided by @MrDoe